OpenStudy (anonymous):

has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (anonymous):

Try solving for x here. \[18 = \frac{x + 25}{2}\]

OpenStudy (anonymous):

1.44???

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

divided 25 by 2 then divided 18 by that number

OpenStudy (anonymous):

Ok. FIrst of all, what is the opposite of division?

OpenStudy (anonymous):

\[18 = \frac{x + 25}{2}\]\[18 \times 2 = \frac{x + 25}{2}\times \frac{2}{1}\]\[36 = \frac{x + 25}{\cancel{2}1} \times \frac{\cancel{2}1}{1}\]\[36 = x + 25\]Can you solve for x now?